Maria Vazquez and La Toya Johnson Plant petunias in a botanical garden. It takes Maria twice as long as Latoya to plant the flowers. Working together, they can plant the flowers in ten hours. How long does it take Latoya to plant the flowers by herself?
Answer by ewatrrr(24785)


Hi,
Let x represent the time it takes for La Toya to plant the garden.
the 2x would represent Maria's time to plant thegarden
Using per hour as an equalizer:
1/x + 1/2x = 1/10hr
Multiplying each term on both sides of the equation by 10x so as all denominatiors = 1
10hr + 5hr = x
15hr = x, the time it takes for La Toya to plant the garden.
checking our answer
1/15 + 1/30 = 2/30 + 1/30 = 3/30 = 1/10